# Self-supervised Brain Lesion Generation for Effective Data Augmentation of Medical Images

### v1

@@ Abstract
Accurate brain lesion delineation is important for planning neurosurgical treatment. Automatic brain lesion segmentation methods based on convolutional neural networks have demonstrated remarkable performance. However, neural network performance is constrained by the lack of large-scale well-annotated training datasets. In this manuscript, we propose a comprehensive framework to efficiently generate new, realistic samples for training a brain lesion segmentation model. We first train a lesion generator, based on an adversarial autoencoder, in a self-supervised manner. Next, we utilize a novel image composition algorithm, Soft Poisson Blending, to seamlessly combine synthetic lesions and brain images to obtain training samples. Finally, to effectively train the brain lesion segmentation model with augmented images we introduce a new prototype consistence regularization to align real and synthetic features. Our framework is validated by extensive experiments on two public brain lesion segmentation datasets: ATLAS v2.0 and Shift MS. Our method outperforms existing brain image data augmentation schemes. For instance, our method improves the Dice from 50.36% to 60.23% compared to the U-Net with conventional data augmentation techniques for the ATLAS v2.0 dataset.
{IEEEkeywords}
Brain Lesion Segmentation, Data Augmentation, Poisson Blending, Prototype Learning, Self-supervised Learning.
@@ 1 Introduction
Data augmentation is a widely used technique to increase training dataset size and diversity in order to improve model performance. Conventional data augmentation strategies include random rotations, brightness adjustment, etc. Although these simple spatial and appearance transformations improve segmentation performance to some extent, they do not fundamentally increase the diversity of the dataset and provide a smaller boost in performance compared to acquiring new data. Recently, methods to perform data augmentation based on image fusion (such as Mixup [3], CutMix [4], etc.) have been developed to increase training dataset diversity. However, these approaches may shift the distribution of the original dataset [5], which is catastrophic for small datasets as the model learns non-representative features in the augmented images [3]. Neural network approaches, such as Generative Adversarial Networks (GANs) [6] have been proposed to synthesize data for model training. However, GAN performance, as with other neural networks, is limited when the training dataset is small creating challenges when generating realistic images with only limited data available. Regardless of the method used to create augmented samples, combining augmented and real samples does not guarantee the segmentation model will learn discriminative features to segment lesions on real samples as there is no supervision in the feature space. Therefore, we raise two questions: How can we generate realistic images that will not shift the original data distribution with limited training samples? How can we effectively use synthetic samples to train a segmentation model to perform well on real samples?
@@
  • We develop a two-stage adversarial autoencoder (AAE) consisting of shape and intensity generation networks. Different from other GANs, we designed a self-supervised approach where we simulate image-label pairs for training our AAE. This enables our AAE to learn a larger distribution of lesions and therefore generate a more realistic range of brain lesions. Furthermore, we use a constrained sampling strategy to guarantee the dataset distribution in the latent space is not shifted.
@@
Figure 1: (a) Real lesions (green arrow) and synthetic lesions (orange arrow) in images from the ATLAS v 2.0 (top) and MS Shift dataset (bottom) demonstrating the synthetic lesions have a similar appearance to real lesions. (b) t-SNE of the image embedding space for real lesions, synthetic lesions, and training samples.
@@ 2.1 Data Augmentation for Data Scarcity
To mitigate the problem of data scarcity when training neural networks, data augmentation techniques are used to increase the training dataset size thereby improving model performance. Conventional data augmentation techniques include shape transformations (flip, rotation, scaling, etc.) and appearance transformations (color jittering, brightness, and contrast adjustment, etc.) [8, 9]. However, the diversity of the dataset achieved by such transformations is limited. Furthermore, the intrinsic characteristics of the training samples are not fundamentally changed, limiting improvements in model performance. The development of GANs enabled more advanced data augmentation where entirely new samples are created by the model. GANs can generate realistic samples for both natural [10] and medical  [11, 12] images. GANs [11, 12] may also be designed to enable image-to-image translation, not generating images from noise, in order to create images of different modalities or characteristics. Nevertheless, GANs require large datasets when training the model to enable accurate image generation. Recently, image-manipulation-based data augmentation techniques [3, 4, 13] have been developed to increase training dataset size and variety by using a set of image manipulation rules to generate new samples. For instance, CutMix [4] cuts and fuses patches from different images to create new training samples. Such techniques must be carefully designed and may shift the distribution of the training dataset especially when the original training dataset size is small [5]. In this study, we combine the strengths of generative models and image-manipulation-based data augmentation techniques to synthesize realistic brain images with limited training samples.
@@ 2.2 Poisson Blending in Deep Learning
Poisson blending is an image processing technique to seamlessly integrate regions of a source image into a target image. In the context of data augmentation, this typically involves identifying the foreground of a source image and integrating this region into a target image. The blending process is guided by the gradient of the source image and the intensity of the target image [7]. Poisson blending generates more coherent images compared to simpler fusion methods such as Copy-Paste [14]. Poisson blending has been used for data augmentation for a variety of natural images [15]. In the context of medical imaging, Tan et al. [16] developed a self-supervised learning strategy to detect abnormalities in chest X-ray images, and abnormalities were generated by fusing image patches into the target image using Poisson blending. Wang et al. [17] utilized Poisson blending to generate retinal images with lesions and different appearances than in the original training dataset to improve the performance of a lesion segmentation model. Lee et al. [18] introduced Poisson blending to the gastric disease classification task for better model generalization ability. However, all of these applications were applied to 2D images, and to our knowledge, Poisson blending has not yet been applied to 3D medical images.
@@ 2.3 Prototype Learning
Prototype networks [19] were first presented for few-shot learning, where the network learns to symbolize each class using a prototype(a.k.a. a representative vector in the embedding space). Each class prototype is typically obtained by computing the average feature of samples belonging to the class. This method was initially applied in image classification where distances between different class prototypes were maximized across training samples [19]. Prototype learning was extended to image segmentation by computing cosine similarity between the class prototypes and individual pixel features. In this context for the test dataset pixels the class with which they have the highest similarity to its prototype. Wang et al. [20] designed a bidirectional prototype alignment mechanism for the few-shot image segmentation task. Kuo et al. [21] utilized class prototypes to augment training samples in the feature space. Xu et al. [22] introduced a cyclic prototype consistency framework for semi-supervised medical image segmentation. In our work, we draw on the idea of prototype consistency to introduce a regularisation term during segmentation model training to align features between synthetic and real samples.
@@ 3 Methodology
Figure 2: Overview of our framework containing three stages. First, the lesion generator is trained via a self-supervised learning strategy and used to generate synthetic lesions based on constrained latent space sampling in Stage I. In Stage II, we seamlessly compose synthetic lesions into full brain images using the proposed Soft Poisson Blending (SPB) to increase the number of training samples. In Stage III, we train the downstream segmentation model with the prototype consistency regularization to align real and synthetic features.
@@ 3.1.1 Shape and Intensity AAE Design
We follow the model architecture proposed in  [23] to design the shape and intensity AAEs. As shown in Fig. 2, each AAE contains an encoder $E$, a decoder $G$, an image discriminator $D_{x}$, and a latent discriminator $D_{z}$. Although the structures are similar, for the intensity AAE we introduce a mask embedding block (MEB) [24] to provide shape guidance when generating the lesion intensity. The MEB is adapted from the SPADE module [25], to use 3D convolutions instead of 2D convolutions.
@@ 3.1.3 Constrained Sampling for New Lesion Synthesis
Figure 3: Lesion mask in three views for (a) real masks and (b) pred-generated masks for the shape AAE model training. Lesion images in three views for (c) real images and (d) pred-generated images for the intensity AAE model training.
@@ 3.2.1 Lesion Location Selection
The brain has a regular anatomical structure, with distinct regions including the ventricle and brain stem. The location of brain lesions depends on the underlying pathology (e.g., stroke and multiple sclerosis). For the datasets in this work, we use FastFurfer [29] to segment the white matter area of the original patient’s brain as a region proposal for lesion location selection. After obtaining the mask for white matter areas, we apply morphological binary erosion operation to shrink the masked region so that the synthetic lesion area will not exceed the mask boundary. We randomly (following a uniform distribution) select a voxel in the masked area as the center of the synthetic lesion.
@@ 3.2.2 Soft Poisson Blending
where $V$ is the guidance vector field and $\nabla$ is the gradient operator. This equation guarantees that: i) The gradient of the foreground content is as close as possible to $V$. ii) The boundary pixel values of the foreground are consistent with the existing $s$, i.e. a seamless transition. The solution under the Dirichlet boundary condition is the Poisson equation:
@@
Figure 4: (a) The target brain image used for the background image. (b) The composited image based on SPB. (c) The composited image based on the original Poisson Blending. (d) The synthetic brain lesion. (e) The guidance vector field used for SPB. (f) The guidance vector field used for the original Poisson Blending. The yellow arrow points to the gradient values on the region boundary.
@@ 4 Experiments
Our framework is evaluated on two public brain segmentation datasets: the ATLAS v2.0 dataset and the Shift MS segmentation dataset as described in Section 4.1. We compare our approach to state-of-the-art (SOTA) methods (Section 4.3) to show the superiority in both lesion synthesis and segmentation tasks. We further conduct ablation studies to validate the effectiveness of each component in our framework (Section 4.4).
4.1 Dataset and Preprocessing
4.1.1 ATLAS v2.0 Dataset
4.1.2 Shift MS Dataset
The Shift MS dataset [32] is a multi-center white matter multiple sclerosis segmentation dataset comprised of i.e. MSSEG-1 [33], ISBI[34], PubMRI [35] and a private dataset collected from the University of Lausanne. We use the three public datasets for model training and evaluation since the private dataset is not publicly available. Detailed information on the dataset is shown in Table 1, the training set (Trn) is used for model training, the in-domain development set (Devin), the out-domain development set (Devout), and the in-domain evaluation set (Evain) are used for evaluating model segmentation performance only.
@@ 4.1.2 Shift MS Dataset
4.2 Implementation Details
The framework is implemented in PyTorch 1.13.1. All model training and validation were performed using an NVIDIA A100 40G GPU. For the lesion generator, the input mask and image size is $64\times 64\times 64$. We used the AdamW optimizer to train the shape and the intensity models, with the learning rate set to $51e-$ 5. The batch size was 4 and the total number of training epochs was 100. For the segmentation model, the input patch size is $128\times 128\times 128$ and the batch size is 2. We used the AdamW optimizer for the segmentation model training with the learning rate set to $31e-$ 3 and consecutively reduced with a cosine annealing strategy. A total of 500 epochs were set for the ATLAS v2.0 dataset and 1000 epochs for the Shift MS dataset. The loss function coefficients are set to $\lambda_{1}=1$, $\lambda_{2}=50$, for each dataset respectively, these values were chosen empirically.
4.3 Model Evaluation
We evaluated our framework on (1) its ability to generate realistic lesions, (2) the performance of the segmentation model trained with images generated using our framework compared to other SOTA models and (3) comparing our framework with other data augmentation techniques for training the segmentation model.
4.3.1 Lesion Synthesis Performance
To evaluate the effectiveness of our generative model and self-supervised training strategy, we compare the synthetic lesions generated by our framework to other generative models. We use the peak signal-to-noise ratio (PSNR) and mean absolute error (MAE) to evaluate lesion realism. The measures are reported in Table 2. For each method, the model is trained either with only real lesions or only pre-generated image-mask pairs created using the self-supervised strategy (see Section 3.1.2). Across all models training with the pre-generated data yields superior metrics compared to the real dataset, underscoring the ability of appropriate synthetic data to enhance model training. Additionally, our approach achieves the highest PSNR and lowest MAE, indicating its robustness and adaptability. Notably, the increased performance of our model is not attributable to the model’s size, as evidenced by the comparison with f-AnoGAN which has the largest number of parameters but the lowest quantitative performance. The performance gains of our method are attributed to its innovative architecture and the utilization of the self-supervised training strategy.
@@ 4.3.1 Lesion Synthesis Performance
AE [36] real 64.11M 31.72 0.0014 30.42 0.0017
synt 64.11M 32.07 0.0011 31.56 0.0014
AAE [37] real 64.41M 30.66 0.0016 31.33 0.0014
synt 64.41M 32.11 0.0010 32.84 0.0009
f-AnoGAN [12] real 78.15M 29.86 0.0019 28.35 0.0023
synt 78.15M 30.22 0.0017 30.05 0.0018
@@
4.3.2 Downstream Segmentation Model Performance
Table 3: Quantitative results of our method compared to other segmentation models for the Shift MS dataset. * indicates that the p-value ¡ 0.05 compared to the second-best model performance computing with a paired Student’s t-test.
@@ 4.3.2 Downstream Segmentation Model Performance
Ours (UNet) $\textbf{78.35$\pm$8.74}^{*}$ 0.91$\pm$1.09 $\textbf{8.15$\pm$12.06}^{*}$ $\textbf{68.52$\pm$15.60}^{*}$ $\textbf{1.73$\pm$3.24}^{*}$ 12.26$\pm$19.37 $\textbf{69.51$\pm$12.63}^{*}$ $\textbf{1.60$\pm$2.17}^{*}$ $\textbf{7.30$\pm$7.75}^{*}$
Ours (Attention UNet) 77.67$\pm$10.98 1.17$\pm$1.53 10.05$\pm$16.44 67.20$\pm$21.04 3.14$\pm$7.59 11.48$\pm$15.63 69.15$\pm$11.59 1.99$\pm$3.81 9.64$\pm$12.31

### v2

@@ Abstract
Accurate brain lesion delineation is important for planning neurosurgical treatment. Automatic brain lesion segmentation methods based on convolutional neural networks have demonstrated remarkable performance. However, neural network performance is constrained by the lack of large-scale well-annotated training datasets. In this manuscript, we propose a comprehensive framework to efficiently generate new samples for training a brain lesion segmentation model. We first train a lesion generator, based on an adversarial autoencoder, in a self-supervised manner. Next, we utilize a novel image composition algorithm, Soft Poisson Blending, to seamlessly combine synthetic lesions and brain images to obtain training samples. Finally, to effectively train the brain lesion segmentation model with augmented images we introduce a new prototype consistence regularization to align real and synthetic features. Our framework is validated by extensive experiments on two public brain lesion segmentation datasets: ATLAS v2.0 and Shift MS. Our method outperforms existing brain image data augmentation schemes. For instance, our method improves the Dice from 50.36% to 60.23% compared to the U-Net with conventional data augmentation techniques for the ATLAS v2.0 dataset.
keywords:
MSC:
@@ 1 Introduction
Data augmentation is a widely used technique to increase training dataset size and diversity in order to improve model performance. Conventional data augmentation strategies include random rotations, brightness adjustment, etc. Although these simple spatial and appearance transformations improve segmentation performance to some extent, they do not fundamentally increase the diversity of the dataset and provide a smaller boost in performance compared to acquiring new data. Recently, methods to perform data augmentation based on image fusion (such as Mixup [43], CutMix [42], etc. ) have been developed to increase training dataset diversity. However, these approaches may shift the distribution of the original dataset [29], which is catastrophic for small datasets as the model learns non-representative features in the augmented images [43]. Neural network approaches, such as Generative Adversarial Networks (GANs) [7] have been proposed to synthesize data for model training. However, the performance of GANs, similar to other neural networks, is constrained by the size of the training dataset, posing challenges in generating realistic images when only limited data is available. Regardless of the method used to create augmented samples, combining augmented and real samples does not guarantee the segmentation model will learn discriminative features to segment lesions on real samples as there is no supervision in the feature space. Therefore, we raise two questions: How can we generate realistic images that will not shift the original data distribution with limited training samples? How can we effectively use synthetic samples to train a segmentation model to perform well on real samples?
@@
  • 1.
    We develop a data augmentation method to enhance the performance of neural networks trained to perform segmentation tasks. It comprises a two-stage adversarial autoencoder (AAE), consisting of shape and intensity generation networks, to generate on-the-fly new foreground regions when training. Distinct from other GANs, we designed a self-supervised approach where we simulate image-label pairs for training our AAE. This enables our AAE to learn a larger distribution of lesions and improve the quality of synthetic results.
  • 2.
@@
Fig. 1: (a) Real lesions (green arrow) and synthetic lesions (orange arrow) in images from the ATLAS v 2.0 (top) and MS Shift dataset (bottom) demonstrating the synthetic lesions have a similar appearance to real lesions. (b) t-SNE of the intensity embedding space for real lesions, synthetic lesions, and training samples. (c) t-SNE of the shape embedding space for real lesions, synthetic lesions, and training samples.
@@ 2.1 Data Augmentation for Data Scarcity
To mitigate the problem of data scarcity when training neural networks, data augmentation techniques are used to increase the training dataset size thereby improving model performance. Conventional data augmentation techniques include shape transformations (flip, rotation, scaling, etc. ) and appearance transformations (color jittering, brightness, and contrast adjustment, etc. [15, 13]. However, the diversity of the dataset achieved by such transformations is limited. Furthermore, the intrinsic characteristics of the training samples are not fundamentally changed, limiting improvements in model performance. The development of GANs enabled more advanced data augmentation where entirely new samples are created by the model. GANs can generate realistic samples for both natural [2] and medical [24, 34] images. GANs may also be designed to enable image-to-image translation, not generating images from noise, in order to create images of different modalities or characteristics. Nevertheless, GANs require large datasets when training the model to enable accurate image generation. Recently, image-manipulation-based data augmentation techniques [43, 42, 44] have been developed to increase training dataset size and variety by using a set of image manipulation rules to generate new samples. For instance, CutMix [42] cuts and fuses patches from different images to create new training samples. Such techniques must be carefully designed and may shift the distribution of the training dataset especially when the original training dataset size is small [29]. In this study, we combine the strengths of generative models and image-manipulation-based data augmentation techniques to synthesize realistic brain images with limited training samples.
@@ 2.2 Poisson Blending in Deep Learning
Poisson blending is an image processing technique to seamlessly integrate regions of a source image into a target image. In the context of data augmentation, this typically involves identifying the foreground of a source image and integrating this region into a target image. The blending process is guided by the gradient of the source image and the intensity of the target image [27]. Poisson blending generates more coherent images compared to simpler fusion methods such as Copy-Paste [6]. Poisson blending has been used for data augmentation for a variety of natural images [20]. In the context of medical imaging, [36] developed a self-supervised learning strategy to detect abnormalities in chest X-ray images, and abnormalities were generated by fusing image patches into the target image using Poisson blending. [38] utilized Poisson blending to generate retinal images with lesions and different appearances than in the original training dataset to improve the performance of a lesion segmentation model. [17] introduced Poisson blending to the gastric disease classification task for better model generalization ability. However, all of these applications were applied to 2D images, and to our knowledge, Poisson blending has not yet been applied to 3D medical images.
@@ 2.3 Prototype Learning
Prototype networks [35] were first presented for few-shot learning, where the network learns to symbolize each class using a prototype(a.k.a. a representative vector in the embedding space). Each class prototype is typically obtained by computing the average feature of samples belonging to the class. This method was initially applied in image classification where distances between different class prototypes were maximized across training samples [35]. Prototype learning was extended to image segmentation by computing cosine similarity between the class prototypes and individual pixel features. In this context for the test dataset pixels the class with which they have the highest similarity to its prototype. [39] designed a bidirectional prototype alignment mechanism for the few-shot image segmentation task. [16] utilized class prototypes to augment training samples in the feature space. [40] introduced a cyclic prototype consistency framework for semi-supervised medical image segmentation. In our work, we draw on the idea of prototype consistency to introduce a regularisation term during segmentation model training to align features between synthetic and real samples.
@@ 3 Methodology
Fig. 2: Overview of our framework containing three stages. First, the lesion generator is trained via a self-supervised learning strategy and used to generate synthetic lesions based on constrained latent space sampling in Stage I. In Stage II, we seamlessly compose synthetic lesions into full brain images using the proposed Soft Poisson Blending (SPB) to increase the number of training samples. In Stage III, we train the downstream segmentation model with the prototype consistency regularization to align real and synthetic features.
@@ 3.1.1 Shape and Intensity AAE Design
We follow the model architecture proposed in  [30] to design the shape and intensity AAEs. As shown in Fig. 2, each AAE contains an encoder $E$, a decoder $G$, an image discriminator $D_{x}$, and a latent discriminator $D_{z}$. Although the structures are similar, for the intensity AAE we introduce a mask embedding block (MEB) [12] to provide shape guidance when generating the lesion intensity. The detailed structure of MEB is shown in Fig. 3. It embeds the mask to the feature space to control the shape of synthetic lesions.
Fig. 3: The detailed structure of MEB block.
@@ 3.1.3 Constrained Sampling for New Lesion Synthesis
Fig. 4: Lesion mask in three views for (a) real masks and (b) pred-generated masks for the shape AAE model training. Lesion images in three views for (c) real images and (d) pred-generated images for the intensity AAE model training.
@@ 3.2.1 Lesion Location Selection
The brain has a regular anatomical structure, with distinct regions including the ventricle and brain stem. The location of brain lesions depends on the underlying pathology (e.g., , stroke and multiple sclerosis). For the datasets in this work, we use FastFurfer [8] to segment the white matter area of the original patient’s brain as a region proposal for lesion location selection. After obtaining the mask for white matter areas, we apply morphological binary erosion operation to shrink the masked region so that the synthetic lesion area will not exceed the mask boundary. We randomly (following a uniform distribution) select a voxel in the masked area as the center of the synthetic lesion.
@@ 3.2.2 Soft Poisson Blending
where $V$ is the guidance vector field and $\nabla$ is the gradient operator. This equation guarantees that: i) The gradient of the foreground content is as close as possible to $V$. ii) The boundary pixel values of the foreground are consistent with the existing $s$, i.e., a seamless transition. The solution under the Dirichlet boundary condition is the Poisson equation:
@@
Fig. 5: (a) The target brain image used for the background image. (b) The composited image based on SPB. (c) The composited image based on the original Poisson Blending. (d) The synthetic brain lesion. (e) The guidance vector field used for SPB. (f) The guidance vector field used for the original Poisson Blending. The yellow arrow points to the gradient values on the region boundary.
@@ 4 Experiments
Our framework is evaluated on two public brain segmentation datasets: the ATLAS v2.0 dataset and the Shift MS segmentation dataset as described in Section 4.1. We compare our approach to existing methods (Section 4.3) to show the superiority in both lesion synthesis and segmentation tasks. We further conduct ablation studies to validate the effectiveness of each component in our framework (Section 4.4).
@@
4.1 Dataset and Preprocessing
4.1.1 ATLAS v2.0 Dataset
4.1.2 Shift MS Dataset
The Shift MS dataset [23] is a multi-center white matter multiple sclerosis segmentation dataset comprised of i.e., MSSEG-1 [5], ISBI[3], PubMRI [18] and a private dataset collected from the University of Lausanne. We use the three public datasets for model training and evaluation since the private dataset is not publicly available. Detailed information on the dataset is shown in Table 1, the training set (Trn) is used for model training, the in-domain development set (Devin), the out-domain development set (Devout), and the in-domain evaluation set (Evain) are used for evaluating model segmentation performance only.
@@ 4.1.2 Shift MS Dataset
AE real 64.11M 31.72 0.0014 30.42 0.0017
[33] synt 64.11M 32.07 0.0011 31.56 0.0014
AAE real 64.41M 30.66 0.0016 31.33 0.0014
[22] synt 64.41M 32.11 0.0010 32.84 0.0009
f-AnoGAN real 78.15M 29.86 0.0019 28.35 0.0023
[34] synt 78.15M 30.22 0.0017 30.05 0.0018
DDPM real 74.23M 32.53 0.0009 32.38 0.0009
[9] synt 74.23M 34.48 0.0007 34.23 0.0008
PNDM real 74.23M 32.56 0.0009 32.32 0.0009
[21] synt 74.23M 34.87 0.0007 34.68 0.0007
@@
Table 3: Quantitative results of our method compared to other segmentation models for the Shift MS dataset. * indicates that the p-value $<$ 0.05 compared to the second-best model performance computing with a paired Student’s t-test.
@@
Ours $\textbf{78.35$\pm$8.74}^{*}$ 0.91$\pm$1.09 $\textbf{8.15$\pm$12.06}^{*}$ $\textbf{68.52$\pm$15.60}^{*}$ $\textbf{1.73$\pm$3.24}^{*}$ 12.26$\pm$19.37 $\textbf{69.51$\pm$12.63}^{*}$ $\textbf{1.60$\pm$2.17}^{*}$ $\textbf{7.30$\pm$7.75}^{*}$
4.2 Implementation Details
The framework is implemented in PyTorch 1.13.1. All model training and validation were performed using an NVIDIA A100 40G GPU. For the lesion generator, the input mask and image size is $64\times 64\times 64$. We used the AdamW optimizer to train the shape and the intensity models, with the learning rate set to $51e-$ 5. The batch size was 4 and the total number of training epochs was 100. For the segmentation model, we use UNet [31] as the backbone model. The input patch size is $128\times 128\times 128$ and the batch size is 2. We used the AdamW optimizer for the segmentation model training with the learning rate set to $31e-$ 3 and consecutively reduced with a cosine annealing strategy. A total of 500 epochs were set for the ATLAS v2.0 dataset and 1000 epochs for the Shift MS dataset. The loss function coefficients are set to $\lambda_{1}=1$, $\lambda_{2}=50$, for each dataset respectively, these values were chosen empirically.
4.3 Model Evaluation
We evaluated our framework on (1) its ability to generate lesions, (2) the performance of the segmentation model trained with images generated using our framework compared to other models and (3) comparing our framework with other data augmentation techniques for training the segmentation model.
4.3.1 Lesion Synthesis Performance
To evaluate the effectiveness of our generative model and self-supervised training strategy, we compare the synthetic lesions generated by our framework to other generative models. We use the peak signal-to-noise ratio (PSNR) and mean absolute error (MAE) to evaluate synthetic results. Structural similarity (SSIM) is not suitable for this scenario because the large proportion of background dominates the small foreground area, leading to an inaccurate representation of the actual image quality. The measures are reported in Table 2. We compare our approach to both GAN and diffusion models. For each method, the model is trained either with only real lesions or only pre-generated image-mask pairs created using the self-supervised strategy (see Section 3.1.2). Across all models training with pre-generated data yields superior metrics compared to the real dataset, underscoring the ability of appropriate synthetic data to enhance model training. Additionally, our approach achieves the highest PSNR and lowest MAE, indicating its robustness and adaptability. Notably, the increased performance of our model is not attributable to the model’s size, as evidenced by the comparison with f-AnoGAN which has the largest number of parameters but the lowest quantitative performance. The performance gains of our method are attributed to its innovative architecture and the utilization of the self-supervised training strategy.
4.3.2 Downstream Segmentation Model Performance